• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 08:39:09 on October 9,2002

    Entry number 87688

    keyword=T3 rates

    In a further investigation of T3 rates, since adaql2 is being rebooted and
    we cannot take data anyway, we ask for the beam to be moved to 0.1 from -0.1 and see the rate drops from ~1.3 kHz to 0.9 kHz.
    But without FFB, the position wanders somewhat anyway.

    We further ask for the raster to be decreased from 2x2 to 1.6x1.6,
    and the rate drops to ~600 Hz.

    This seems to indicate most of the T3 rate really is neutrals from
    the radiator.

    MCC indicates they can put the Compton locks on, unrelated
    to the lack of FFB, and this should stabilize our beam somewhat.